OpenStudy (anonymous):

the integral of ((e^x +1)^(-1/2))e^x dx from ln8 to ln 24

OpenStudy (anonymous):

\[ u= 1 + e^x \]

OpenStudy (anonymous):

du = e^x dx

OpenStudy (anonymous):

\[ \int ((e^x +1)^{-1/2})e^x dx=\int u^{-1/2} du \]

OpenStudy (anonymous):

Do you know how to finish it from here

OpenStudy (anonymous):

No, are the limits still ln8 to ln24 or do they change

OpenStudy (anonymous):

Get the integral in terms of u, replace u by its value and keep the same limits

OpenStudy (anonymous):

Or you can change the limit before the integration.

OpenStudy (anonymous):

So is the final answer going to be 4?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Thanks